OpenStudy (andriod09):

how do you solve : \[\frac{7}{4}c-\frac{8}{7}=\frac{13}{4}+\frac{2}{7}c\]

OpenStudy (aravindg):

bring 2/7 c to left hand side

OpenStudy (aravindg):

8/7 to right hand side

OpenStudy (aravindg):

Then you can add like terms

OpenStudy (tkhunny):

Various ways to go about it. If it were me, I'd multiply the world by 28.

ganeshie8 (ganeshie8):

multiply the whole equation with 7*4

OpenStudy (andriod09):

Concidering that @AravindG was helping me, i'll go with his help. Thanks to you guys though. :P

OpenStudy (aravindg):

@andriod09 multiplying by 7*4 is better a first step

OpenStudy (aravindg):

It makes things easier

OpenStudy (anonymous):

\[\frac{ 7 }{ 4 }c - \frac{ 2 }{ 7 }c=\frac{ 13 }{ 4 }+\frac{ 8 }{ 7 }\] \[\frac{ (49-8) }{ 28 }c=\frac{ 91+32 }{ 28 }\] Multiply by 28 on both sides \[(49-8)c=91+32\] \[41c=123\] \[c=\frac{ 123 }{ 41 }\] c=3

OpenStudy (aravindg):

omg ^ :)

OpenStudy (aravindg):

Any doubts?

OpenStudy (andriod09):

Nope. Thanks guys. I hate fractions, i suck at them, and never get them right.

OpenStudy (aravindg):

You see when you multiply by 7*4 you get rid of the fraction ..so it becomes simpler

OpenStudy (aravindg):

The qn simply reduces to 49c -32=91+8c

OpenStudy (aravindg):

gt it?